This Question: 2 pts 21 of 35 (2 complete) The reason investors buy bonds is to O A. own controlling interest in the company O B. receive dividend payments O C. earn interest D. exercise voting rights in a company
media%2Fbae%2Fbae3967b-02e2-4698-a28e-04
media%2F9a3%2F9a364c04-3810-4a1d-ae34-23
media%2Fa6f%2Fa6f8fc18-d5f2-47d0-bfd0-cc
media%2F836%2F836e7916-6d39-40d4-9071-e5
Answer #1

Answers

  • Question 21

Correct answer is Option ‘C’ Earn Interest. Interest is paid on Bonds Payable. No dividend, voting rights or controlling interest is provided under Bonds Issue.

  • Question 22

Discount = 35000 – 15700 = $ 19,300

No. of interest payment = 6 years x 2 = 12

Discount to be amortised with each interest payment = 19300 / 12 = $ 1608.33

Cash Interest paid on June 30 = 35000 x 14% x 6/12 = $ 2450

Interest expense to be debited by = 2450 + 1608.33 = $ 4,058 = Option ‘A’

  • Question 23

Adjustment to Net Income:

Add: Decrease in Accounts receivables = $ 3000
Less: Increase in Inventory = $ 29000
Add: Depreciation expense = $ 7000

Adjustments = 3000 + 7000 – 29000 = $ (19,000) = Option ‘B’

  • Question 24

Working capital = Current Assets – Current Liabilities.

= 42500 + 27000 + 94000 + 156000 – 270000 = $ 49,500 = Option ‘A’

  • Question 25

Increase in Current Assets is deducted from Net Income under Operating activities section.

Since Inventories are part of Current Assets and has INCREASED, correct answer is Option ‘B’: Subtraction from Net Income under the operating activities section.